Standard Product Costs
Deerfield Company manufactures product M in its factory. Production of M requires 2 pounds of material P, costing $8 per pound and 0.5 hour of direct labor costing, $14 per hour. The variable overhead rate is $12 per direct labor hour, and the fixed overhead rate is $16 per direct labor hour. What is the standard product cost for product M?

Direct material Answer
Direct labor Answer
Variable overhead Answer
Fixed overhead Answer
Standard product cost per unit Answer
Answer #1

А 1 Direct material 2 Direct labor 3 Variable overhead 4 Fixed overhead 5 Standard product cost per unit a 00 $

А B 1 Direct material =2*8 2 Direct labor =0.5*14 3 Variable overhead =0.5*12 4 Fixed overhead =0.5*16 5 Standard product cos